# A presentation of the pure cactus group of degree four

### v1

@@ 3. Preliminary
Here $H_{g}(e)$ denotes the half space defined by $\{x\in\mathcal{C}\mid d(x,e)<d(x,g\cdot e)\}$ for $g$ in $PJ_{4}$.
[…]
Let $g$ be a non-trivial element of $PJ_{4}$. By Lemma 4.3, if $g$ satisfies that $d(e,g\cdot e)$ is $4$ or less, then it is equal to $4$ and $g=g_{i}^{\pm 1}$ for some $i$. Therefore, $H_{g}(e)$ is equal to $H_{g_{i}^{\pm 1}}(e)$ for some $i$ when $d(e,g\cdot e)$ is 4 or less.
If $d(e,g\cdot e)$ is $6$ or more, we obtain the following.
$\{x\mid d(e,x)<3\}\subset H_{g}(e)$
[…]
It remains to show that $d(e,g\cdot e)\neq 5$ on $\mathcal{C}$ for any $g$ in $PJ_{4}$.
Recall that the image of the standard generators $s_{12}$, $s_{23}$, $s_{34}$, $s_{13}$, $s_{24}$ of $J_{4}^{\prime}$ under $\pi$ are the transpositions $(12)$, $(23)$, $(34)$, $(13)$, $(24)$ in $S_{4}$, and also that $\pi(s_{14})=(14)(23)$. See Section 2.
Let $g$ be an element in $PJ_{4}$. It means that $\pi(g)=e$ in $S_{4}$.
If $g$ is contained in $J_{4}^{\prime}$, then $g\cdot e=g$ is in $J_{4}^{\prime}$. Thus $d(e,g\cdot e)=d(e,g)=5$ implies $g$ has word length 5 with respect to the generators $s_{12}$, $s_{23}$, $s_{34}$, $s_{13}$, $s_{24}$ in $J_{4}^{\prime}$. It would imply that $\pi(g)$ is an odd permutation, contradicting that $\pi(g)=e$.
If $g$ is not contained in $J_{4}^{\prime}$, then $g\cdot e=gs_{14}$ is in $J_{4}^{\prime}$ by definition of the action $\Gamma$. Thus $d(e,g\cdot e)=d(e,gs_{14})=5$ implies $gs_{14}$ has word length 5 with respect to the generators $s_{12}$, $s_{23}$, $s_{34}$, $s_{13}$, $s_{24}$ in $J_{4}^{\prime}$, that is, $\pi(gs_{14})$ must be an odd permutation. Since $\pi(s_{14})=(14)(23)$, it would imply that $\pi(g)$ is an odd permutation, contradicting that $\pi(g)=e$.
As a result, $D(e)$ contains the interior of $\widetilde{D}$, and it concludes that the polygon $\widetilde{D}$ is the Dirichlet polygon with center $e$ for $PJ_{4}$ as a subgroup of $\mathrm{Isom}\;\mathcal{C}$. ∎
@@
By Lemma 4.5, the interior of $\widetilde{D}$ is the Dirichlet domain for the action $\Gamma$ by $PJ_{4}$. Since the group which acts on a topological space freely and cocompactly is discontinuous, togethter with Proposition 4.2, $PJ_{4}$ is shown to be discontinuous. By Proposition 4.1, $\mathcal{C}$ is geodesically conected, geodesically complete, and finitely compact. Thus, applying Proposition 3.1 to the interior of $\widetilde{D}$, we obtain that $\widetilde{D}$ is a locally finite fundamental polygon for $PJ_{4}$. By Proposition 3.2, $PJ_{4}$ is generated by the elements $g_{1},\dots,g_{10}$ and their inverses.
Next, we will obtain the relations with respect to $g_{1},...,g_{10}$ by applying Proposition 3.3. The generators $g_{1},...,g_{10}$ of $PJ_{4}$ give identifications on (the sides of) the polygon $\widetilde{D}$. For example, by
@@
We remark that, by calculating the Euler characteristic, we see that the surface obtained as $\widetilde{D}^{*}$ is homeomorphic to the connected sum of the five projective plane.
@@ 5. Proof of corollaries
By virtue of geometric descriptions, we could find the correspondence between $\alpha_{i}$ and $g_{j}$ in the proof above as follows. The painted regions in Figure 5 give five Möbius bands embedded in $\mathcal{C}/PJ_{4}$. After removing them, we can rebuild the remaining parts to obtain a sphere with five holes. See Figure 6. Then we can confirm that $\mathcal{C}/PJ_{4}$ is homeomorphic to the connected sum #$\#_{5}\mathbb{RP}^{2}$ of five real projective planes. Setting the closed curve $\alpha_{2}$ in #$\#_{5}\mathbb{RP}^{2}$ shown in Figure 6, we obtain a correspondence $\alpha_{2}\longmapsto g_{8}^{-1}g_{4}g_{9}g_{2}g_{10}\in PJ_{4}$ by considering a lift of $\alpha_{2}$ in $\mathcal{C}$. In the same way, we can obtain the corresponding elements in $PJ_{4}$ for $\alpha_{1}$, $\alpha_{3}$, $\alpha_{4}$, $\alpha_{5}$.

### v2

@@ 4. Key propositions
Here, $H_{g}(e)$ denotes the half-space defined by $\{x\in\mathcal{G}\mid d(x,e)<d(x,g\cdot e)\}$ for $g$ in $PJ_{4}$.
We prove that $\widetilde{D}_{1}$ coincides with $D(e)$. By the definition of the Dirichlet domain, $\widetilde{D}_{1}$ contains $D(e)$. Thus, it suffices to show the converse, i.e., that $H_{g}(e)$ contains $\widetilde{D}_{1}$ for any nontrivial element $g$ in $PJ_{4}$.
Let $g$ be a nontrivial element of $PJ_{4}$. By Lemma 5.1, if $g$ satisfies $d(e,g\cdot e)\leq 4$, then $d(e,g\cdot e)=4$ and $g=g_{i}^{\pm 1}$ for some $i$. Therefore, $H_{g}(e)$ is equal to $H_{g_{i}^{\pm 1}}(e)$ for some $i$ when $d(e,g\cdot e)\leq 4$.
If $d(e,g\cdot e)\geq 6$, we obtain the following:
$\{x\mid d(e,x)<3\}\subset H_{g}(e).$
Since the word lengths of the vertices in $\widetilde{D}_{1}$ are less than $3$, we have $d(e,x)<3$ for any $x\in\widetilde{D}_{1}$. Thus,
$\bigcap\{H_{g}(e)\mid d(e,g\cdot e)\geq 6\}\supset\widetilde{D}_{1}.$
It remains to show that $d(e,g\cdot e)\neq 5$ on $\mathcal{G}$ for any $g\in PJ_{4}$. Recall that the images of the standard generators $s_{12}$, $s_{23}$, $s_{34}$, $s_{13}$, $s_{24}$ of $J_{4}^{\prime}$ under $\pi$ are the transpositions $(12)$, $(23)$, $(34)$, $(13)$, $(24)$ in $S_{4}$, and also that $\pi(s_{14})=(14)(23)$. See Section 2.
Let $g$ be an element in $PJ_{4}$. This means that $\pi(g)=e$ in $S_{4}$.
If $g$ is contained in $J_{4}^{\prime}$, then $g\cdot e=g$ is in $J_{4}^{\prime}$. Thus $d(e,g\cdot e)=d(e,g)=5$ implies that $g$ has word length 5 with respect to the generators $s_{12}$, $s_{23}$, $s_{34}$, $s_{13}$, $s_{24}$ in $J_{4}^{\prime}$. This would imply that $\pi(g)$ is an odd permutation, contradicting $\pi(g)=e$.
If $g$ is not contained in $J_{4}^{\prime}$, then $g\cdot e=gs_{14}$ is in $J_{4}^{\prime}$ by definition of the action $\Gamma$. Thus $d(e,g\cdot e)=d(e,gs_{14})=5$ implies $gs_{14}$ has word length 5 with respect to the generators $s_{12}$, $s_{23}$, $s_{34}$, $s_{13}$, $s_{24}$ in $J_{4}^{\prime}$, and $\pi(gs_{14})$ must be an odd permutation. Since $\pi(s_{14})=(14)(23)$, it implies that $\pi(g)$ is an odd permutation, contradicting $\pi(g)=e$.
As a result, $D(e)$ contains the interior of $\widetilde{D}_{1}$. It concludes that $\widetilde{D}_{1}=\mathrm{int}(\widetilde{D}\cap\mathcal{G})$ is the Dirichlet domain with center $e$ on $\mathcal{G}$ with respect to the action $\Gamma_{1}$ of $PJ_{4}$. ∎
@@
By Lemma 5.3, it follows that the interior of $\widetilde{D}$ is the Dirichlet domain on $\mathcal{C}$, and thus $\widetilde{D}$ is a Dirichlet polygon on $\mathcal{C}$ with respect to the action $\Gamma$ of $PJ_{4}$ on $\mathcal{C}$. This is because the action $\Gamma$ on $\mathcal{C}$ is a natural extension of $\Gamma_{1}$ on $\mathcal{G}=\mathcal{C}^{(1)}$, which is an isometric action on $\mathcal{C}$, and $\widetilde{D}$ is the convex hull in $\mathcal{C}$ of $\widetilde{D}\cap\mathcal{G}$.
Since a group acting freely and cocompactly on a topological space is discontinuous, together with Proposition 4.2, it follows that $PJ_{4}$ is discontinuous. By Proposition 4.1, $\mathcal{C}$ is geodesically connected, geodesically complete, and finitely compact. Thus, applying Proposition 3.1 to the interior of $\widetilde{D}$, we obtain that $\widetilde{D}$ is a locally finite fundamental polygon for $PJ_{4}$. By Proposition 3.2, $PJ_{4}$ is generated by the elements $g_{1},\dots,g_{10}$ and their inverses.
Next, we will obtain the relations for $g_{1},\dots,g_{10}$ by applying Proposition 3.3. The generators $g_{1},\dots,g_{10}$ of $PJ_{4}$ define identifications on the sides of the polygon $\widetilde{D}$. For example, by
@@
We remark that, by calculating the Euler characteristic, we see that the surface obtained from $\widetilde{D}$ by gluing its sides via $g_{j}$’s is homeomorphic to the connected sum of the five projective plane.
@@ 6. Proof of corollaries
By virtue of geometric descriptions, we could find the correspondence between $\alpha_{i}$ and $g_{j}$ in the proof above as follows. The painted regions in Figure 5 represent five Möbius bands embedded in $\mathcal{C}/PJ_{4}$. After removing them, the remaining parts can be reassembled to form a sphere with five holes. See Figure 6. Then, we can confirm that $\mathcal{C}/PJ_{4}$ is homeomorphic to the connected sum #$\#_{5}\mathbb{RP}^{2}$ of five real projective planes. By setting the closed curve $\alpha_{2}$ in #$\#_{5}\mathbb{RP}^{2}$ shown in Figure 6, we obtain a correspondence
$\alpha_{2}\longmapsto g_{8}^{-1}g_{4}g_{9}g_{2}g_{10}\in PJ_{4}$
by considering a lift of $\alpha_{2}$ in $\mathcal{C}$. Similarly, we obtain the corresponding elements in $PJ_{4}$ for $\alpha_{1}$, $\alpha_{3}$, $\alpha_{4}$, and $\alpha_{5}$.